OpenStudy (anonymous):

Which property of equality can be used to solve the equation n plus nine equals negative three? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (mathstudent55):

You have n + 9 = -3 What do you need to do to the 9 on the left side to get rid of it from there, so you end up with n alone?

OpenStudy (anonymous):

subtraction?

OpenStudy (mathstudent55):

Right. If you subtract 9 from both sides, you end up with n by itself on the left side and -12 on the right side: n + 9 = -3 -9 -9 ----------- n = -12 n = -12 So which property was used?